OpenStudy (anonymous):

The numbers 45, 28, and 53 form a Pythagorean triple. true of false

OpenStudy (anonymous):

Square all three, if the smaller two add up to the larger one, then true, otherwise false.

OpenStudy (anonymous):

So that's 45^2+28^2=?53^2

OpenStudy (anonymous):

either \[45^2+28^2=53^2\] in which case the answer is "yes" or \[45^2+28^2\neq 53^2\] in which case the answer is "no"